This is the question. Nathan's grandpa is 6 times as old as Nathan was 2 years ago. If Nathan's grandpa is 54, how old in Nathan?
Answer by lwsshak3(11628)

This is the question. Nathan's grandpa is 6 times as old as Nathan was 2 years ago. If Nathan's grandpa is 54, how old in Nathan?
**
let x=Nathan's present age
x-2=Nathan's age 2 yrs ago
grandpa is 6 times as old as Nathan was 2 years ago.
54=6(x-2)=6x-12
6x=66
x=11
Nathan's present age=11